Please help me solve!! than you!

Question 7 During 2013, Grouper Corporation spent $152,640 in research and development costs. As a result, a new product called the New Age Piano was patented. The patent was obtained on October 1, 2013, and had a legal life of 20 years and a useful life of 10 years. Legal costs of $28,080 related to the patent were incurred as of October 1, 2013
Your answer is partially correct. Try again. On June 1, 2015, Grouper spent $17,280 to successfully prosecute a patent infringement suit. As a result, the estimate of useful life was extended to 12 year from June 1, 2015. Prepare all journal entries required in 2015 and 2016. (Round answers to O decimal places, e.g. 8,564. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts Date Account Titles and Explanation Debit Credit 2015 Patents 17280 17280 Cash (To record patents) 2015 Amortization Expense 840 Patents 840 To record amortization expense) 2016 Amortization Expense 1440 Patents 1440
Your answer is partially correct. Try again. In 2017, Grouper determined that a competitor's product would make the New Age Piano obsolete and the patent worthless by December 31, 2018. Prepare all journal entries required in 2017 and 2018. (Round answers to 0 decimal places, e.g. 8,564. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter O for the amounts.) Debit Account Titles and Explanation Credit Amortization Expense 7500 Patents 7500
Answer #1

Solution

Amortization expense in 2013[(28080÷10)×3/12]=$702

Amortization expense in 2014(28080÷10) =$2808

Amortization expense in 2015:

Date Account titles and explanation Debit Credit
2015 Amortization expense $3148
Patents(1170+1978) $3148

(To record amortization expense) (Jan 1 -June 1: (28080÷10) ×5/12=$1170)

{june 1-Dec 31: [(28080-702-2808-1170+17280) =$40680]; [(40680÷12) ×7/12=$1978}

2016 Amortization expense $3390
Patents(40680/12) $3390
(To record Amortization expense)

In 2017 and 2018:

Date Account titles and explanation Debit Credit
2017 and 2018 Amortization expense 17656
Patents(35312/2) 17656
(To record amortization expense) (40680-1978-3390=$35312)